A parallel conductor transmission line #6 with a distance of 30.48 cm (12 in.)

Made of AWG copper wire (diameter =4.11mm (0.162 inch), sigmax= 58 MS/m). Conductors

Assume that there is air between Neglecting the internal inductance; L, C. G's per meter

their values; Find the dc resistance and the ac resistance at 1 MHz


1
Expert's answer
2021-06-23T05:54:19-0400

Inductance (L) ="2*10^{-7} \ln\frac{d}{r'}=2*10^{-7} \ln\frac{30.48*10^{-2}}{1.6*10^{-3}}=10.49* 10^{-7}H/m"

Capacitor (C) ="\frac{2 \pi \epsilon_0}{ \ln (\frac{d}{r})}=\frac{2 \pi *8.85*10^{-12}}{ \ln (\frac{30.48*10^{-2}}{4.11/2 *10^{-3}})}=11.122*10^{-12}F/m"

DC resistance, "R= \rho \frac{l}{A}= \frac{1}{\delta}\frac{l}{A}= \frac{1}{58*10^6}\frac{30.48*10^{-2}}{13.267*10^{-6}}=39.63* 10^{-3} \Omega"

"G= \frac{1}{R}=25.23 \Omega ^{-1}"

Approximate AC resistance

"Z=R_{AC}=R+j wL=39.63*10^{-3}+j2 \pi*10^6*10.49*10^{-7}=39.63*10^{-2}+j6.6"